2.) A sample of neon gas collected at a pressure of 946 mm Hg and a temperature of 284 K has a mass of 20.6 grams. The volume of the sample is____ L.

3.) What volume of hydrogen gas is produced when 28.8 g of sodium reacts completely according to the following reaction at 25 °C and 1 atm?

sodium (s) + water(l)sodium hydroxide (aq) + hydrogen(g)

______liters hydrogen gas
1
Expert's answer
2020-10-19T14:04:08-0400

2)at stp volume of a given mass of gas is 22.4L when pressure is 760mmHg and temperature 273k

Since, P1V1/T1 = P2V2/T2 and; P1= 760mmHg, V1=22.4L, T1= 273K , P2= 946mmHg, V2 = n T2 =284K

To get n (V2) = 760.ommHg*22.4L/273.0K= 946.0mmHg*n/284K

n =760*22.4*284/273*946

n= 18.7L

Therefore the volume of the sample of neon is 18.7L


3)Given that 28.8g of Na reacts with water at r.t.p( pressure 1atm and temperature 25°C)

Equation: 2Na(s) + 2H2O(l) = 2NaOH(aq) + H2(l)

moles of Na = 28.8g/23.0g/m = 1.252 moles

Mole ration of Na to H2 = 2:1

Moles of H2 = 1*1.252/2= 0.621 moles

if volume of 1mol of H2 at r.t.p is 24.0L;

0.621moles = 24.0L*0.621moles/1 mol= 15.026L

Volume of H2 is therefore 15.026L